OpenStudy (anonymous):

This problem is driving me crazy. A box is given a push so that it slides across the floor. How far will it go, given that the coefficient of kinetic friction is 0.15 and the push imparts an initial speed of 3.5m/s? The answer is 4.2m but I don't know how to get it.

OpenStudy (anonymous):

d=v^2/2ug d=(3.5)^2/(0.15)(9.8m/s^2)

OpenStudy (anonymous):

this should give approx. 4.2m

OpenStudy (anonymous):

How did you get \[x -x_{0} = \frac{v^{2}}{2\mu g}\]?

OpenStudy (anonymous):

Oh wait, I see how. Thanks!

OpenStudy (anonymous):

ok so F=umg work done by frictional force is W=umgd andwe know that KE=1/2mv^2 1/2mv^2=umgd

OpenStudy (anonymous):

:)